Shift report 20020119 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat Jan 19 2002 - 08:11:14 EST

  • Next message: Apache: "Shift report 20020119 08:00-16:00"

    ------------------------------------------------------------
    Date: Sat Jan 19 08:11:14 2002, Shift: 00:00-08:00
    Supervisor: RD / BHS, with: 
    Log book#12 pages 86 - 88
    
    Shift Summary:
    ------------------------------------------------------------
    Two stores tonight, one ending at 03:15 and a second lasting
    from 06:05 to 07:30. MCR reported polarizing problems, and
    rates ranged between low and very low.
    
    FFS was moved to 12deg, 1/2 A, before I came
    BFS is still at 3 deg - are we supposed to move it? I tried to align its two parts a bit better, but it would not move. Since the data rate was so low I decided not to wake anybody to ask...
    MRS is still at 90deg, 6B
    
    Data runs:
    6676 - 7k
    6677 - 4k
    
    Pedestal runs:
    6678 - TPC ped., installed
    6679 - DVM ped., running at end of shift
    
    T3 and T5 had a trip each at the start of the second store, 
    but they both reset nicely.
    
    And of course: My humblest and deepset apologies to Roman who was not relieved on time and Ramiro who was subsequently dragged out of bed becuse a short power-outage in Fleming House had 'tripped' my alarm-clock... Note to self: Get battery-driven clock!
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sat Jan 19 2002 - 08:12:00 EST